OpenStudy (anonymous):

What is a counterexample for the conjecture? Conjecture: The product of two positive numbers is greater than the sum of tge two numbers. A.) 3 and 5 B.) 2 and 2 C.) A counterexample exists, but it is not shown above. D.) There is no counterexample. The conjecture is true.

OpenStudy (tkhunny):

2+2 = ?? 2*2 = ?? 1+1 = ?? 1*1 = ?? Keep it all above 2 and the conjecture is good.

OpenStudy (anonymous):

What?

OpenStudy (tkhunny):

Answer the question marks. You should see it.